# SkyOctopus: Enabling Low-Latency Mobile Satellite Network through Multiple Anchors

### v1

@@ I Introduction
Nowadays, we are witnessing the rapid development of low earth orbit (LEO) satellite constellations, such as SpaceX’s Starlink [1, 2, 3, 4, 5], Amazon’s Kuiper [6], and OneWeb [7]. These constellations, through their dense satellite distribution and use of inter-satellite links (ISLs), are capable of providing global internet and communication services. They can complement terrestrial mobile networks in underserved areas in a cost-effective manner. Meanwhile, as the developer of 5G, 3GPP has explicitly stated that non-terrestrial networks (NTN) will be an essential component of future 5G-Advanced and 6G networks [8, 9, 10, 11, 12]. In practice, it has become a trend for telecom operators and satellite network providers to collaborate to build mobile satellite networks, as exemplified by partnerships such as T-Mobile with SpaceX [13] and AT&T with AST [14].
@@ IV Architecture Design
In this section, we provide a detailed explanation of our proposed SkyOctopus architecture. First, we provide a comprehensive overview of SkyOctopus, along with detailed discussions on the modifications to network function, signaling, and protocol. Then, we propose a fine-grained anchor point selection strategy, which includes initial location-based anchor point selection and the path update mechanism driven by both computation and measurement. Finally, we discuss the anchor distribution problem and propose an algorithm to minimize latency by optimally selecting deployment locations for a fixed number of anchor points.
@@ IV-B3 Working Stages
Simultaneously, the path update mechanism is triggered. First, the S-UPF sends the IP address as a signaling parameter to the core network, which instructs anchor points to send probe messages to detect their latency to the target IP. Based on the output of the computational module and the latency returned by the anchor points, the core network calculates the latency from the S-UPF to the target server via each anchor point. The core network then creates a PDR for the S-UPF, indicating the optimal anchor for the target IP with a higher matching priority than the initially built PDR. The path update mechanism remains active, and if the latency through an alternative anchor point is superior to the current one, the core network initiates a PDR modification signaling instead of creation.
@@ V-A Experimental Setup
Ground Station and User Traffic: We use 40 servers from AWS’s 29 available zones [30] as ground stations. Among these, 20 ground stations will be selected as anchor points, which means $h=20$.
For user traffic, we randomly generate 100 users navigating through different locations in the Atlantic Ocean. Additionally, we select 25 target servers randomly from the top 50 most visited websites globally [31].
@@ VII Discussion and limitation
VII-A Content Delivery Network
In this paper, we consider that users always access the specific target server. However, the content delivery network (CDN) allows users to be served by the nearest available server. In such cases, our ICMP-based probing of the target IP might not reflect the actual network conditions—a closer server could potentially serve users. This situation can result in SkyOctopus selecting suboptimal anchor points and paths, thereby affecting its performance in reducing user latency to some extent. We believe that cooperation among anchor points can better adapt to CDN and leave it for future work.
VII-B Satellite Computing
@@ VIII Conclusion
Circuitous routing is a significant challenge faced by mobile satellite networks, which is largely caused by the design of a single-anchor session. In this paper, we propose an advanced mobile satellite network architecture. We introduce SkyOctopus, which avoids circuitous routing through multi-anchor connections and achieves low-latency access for users with a fine-grained anchor point selection strategy. Furthermore, we construct a prototype of SkyOctopus using enhanced Open5GS and UERANSIM, and conduct experiments to evaluate its performance. The results show that our solution can reduce end-to-end latency by up to 53% and session establishment time by 86%. As a potential future direction, we are looking forward to extending our method to improve the performance of various applications such as large language models [47, 48, 49, 50, 51], deep learning system [52, 53, 54, 55, 56, 57, 58, 59, 60, 61] in satellite networks.

### v2

@@ I Introduction
Nowadays, we are witnessing the rapid development of low earth orbit (LEO) satellite constellations, such as SpaceX’s Starlink [1], Amazon’s Kuiper [2], and OneWeb [3]. These satellite constellations, with their dense satellite distribution and inter-satellite links (ISLs), provide global internet and communication services while complementing terrestrial mobile networks in underserved areas in a cost-effective manner.
Meanwhile, as the developer of 5G, 3GPP has explicitly stated that non-terrestrial networks (NTN) will be an essential component of future 5G-Advanced and 6G networks [4, 5, 6]. In practice, it has become a trend for telecom operators and satellite network providers to collaborate to build mobile satellite networks, as exemplified by partnerships such as T-Mobile with SpaceX [7] and AT&T with AST [8].
@@ IV Architecture Design
In this section, we provide a detailed explanation of our proposed SkyOctopus architecture. First, we provide a comprehensive overview of SkyOctopus, along with detailed discussions on the modifications to network function, signaling, and protocol.
Then, we propose a fine-grained anchor point selection strategy, which includes initial location-based anchor point selection and the path update mechanism driven by both computation and measurement.
Finally, we discuss the anchor distribution problem and propose an algorithm to minimize latency by optimally selecting deployment locations for a fixed number of anchor points.
@@ IV-B3 Working Stages
Simultaneously, the path update mechanism is triggered. First, the S-UPF sends the IP address as a signaling parameter to the core network, which instructs anchor points to send probe messages to detect their latency to the target IP. Based on the output of the computational module and the latency returned by the anchor points, the core network calculates the latency from the S-UPF to the target server via each anchor point. The core network then creates a PDR for the S-UPF, indicating the optimal anchor for the target IP with a higher matching priority than the initially built PDR. The path update mechanism remains active, and if the latency through an alternative anchor point is superior to the current one, the core network initiates a PDR modification instead of creation.
@@ V-A Experimental Setup
Ground Station and User Traffic: We use 40 servers from AWS’s 29 available zones [24] as ground stations. Among these, 20 ground stations will be selected as anchor points, which means $h=20$. For user traffic, we randomly generate 100 users navigating through different locations in the Atlantic Ocean. Additionally, we select 25 target servers randomly from the top 50 most visited websites globally [25].
@@ VII Discussion and limitation
VII-A Satellite Computing
@@
VII-B Content Delivery Network
In this paper, we consider that users always access the specific target server. However, the content delivery network (CDN) allows users to be served by the nearest available server. In such cases, SkyOctopus still provides users with the optimal anchor point selection, but the services requested by users may cluster around a specific anchor point, leading to an increase in its utilization. We believe that cooperation among anchor points can better adapt to CDN and leave it for future work.
@@ VIII Conclusion
Circuitous routing is a significant challenge faced by mobile satellite networks, which is largely caused by the design of a single-anchor session. In this paper, we propose an advanced mobile satellite network architecture. We introduce SkyOctopus, which avoids circuitous routing through multi-anchor connections and achieves low-latency access for users with a fine-grained anchor point selection strategy. Furthermore, we construct a prototype of SkyOctopus and conduct experiments to evaluate its performance. The results show that our solution can reduce end-to-end latency by up to 53% and session establishment time by 86%.